At what Temperature is the speed of sound in methane, CH4 equal to the speed of the sound in air at room temperature?


1
Expert's answer
2021-04-13T06:28:06-0400

Speed of sound in air at room temperature is 346 m/s. For methane, it can be expressed as


"c=\sqrt{\frac{\gamma RT}{M}},\\\space\\
T=\frac{c^2M}{\gamma R}=\frac{346^2·0.016}{1.313·8.314}=175\text{ K}."